OpenStudy (anonymous):

(x-1)+(x+18)+(x+10)+3x+2x+x=360 find x

OpenStudy (anonymous):

(x-1)+(x+18)+(x+10)+3x+2x+x=360 x-1+x+18+x+10+3x+2x+x = 360 (x+x+x+3x+2x + x) +(-1+18+10) = 360 9x+27 = 360 9x = 360-27 9x = 333 x = 37